# On the cardinality of measures of maximal relative entropy for smooth skew products

### v1

@@ 4. Measures of Relative Maximal Entropy on Topological Markov Shifts
We say that $\mu_{1}$ and $\mu_{2}$ are $\nu$-relative orthogonal if
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, if $\mu_{1}$ and $\mu_{2}$ are ergodic measures of $\nu$-maximal relative entropy then $\mu_{1}$ and $\mu_{2}$ are $\nu$-relative orthogonal.
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, $(X,\sigma_{X})$ can only admit countably many ergodic measures of $\nu$-maximal relative entropy.
@@
Given an invariant measure $\mu$ of $(X,\sigma)$ such that $\pi_{*}\mu=\nu$, observe that $\mu_{n}:=(\mathrm{proj}_{X_{n}})_{*}\mu$ satisfies $(\pi_{n})_{*}\mu_{n}=\nu$. From the Kolmogorov-Sinai theorem (see [46, Theorem 4.9]) we have that
@@
where $\beta_{n}$ is the partition of $X_{n}$ 11-cylinders at coordinate 0. Since $\mathcal{B}(X)=\mathcal{B}\left(\bigcup_{n}\sigma(\alpha_{n})\right)$ from [46, Theorem 4.14]
@@ 4.1. Petersen-Quas-Shin construction
Now, we quote a result from [39] which is contained in the proof of Theorem 1 (see [39, Equation (45)]).
@@
We assume the notations of Theorem 4.6 and Lemma 4.8. Assume, for contradiction, that the theorem is false. Then there exist two ergodic measures of $\nu$-maximal relative entropy, $\mu^{1}$ and $\mu^{2}$, such that
@@
Define $\mu^{3}=(\mathrm{proj}^{3}_{X_{n}})_{*}\eta$ by the construction at the beginning of Section 4.1 (we keep the same notation, omitting the index $n$). We will show that
@@
$h_{\mu^{i}}(X)=\lim_{n\to\infty}h_{\mu^{i}_{n}}(X_{n})\ \text{for }i\in\{1,2,3\},\ \text{where}\ \mu_{n}^{i}:=(\pi_{n})_{*}\mu^{i}.$
@@
on an $\eta$-positive measure set contained in $S:=\{\,z\in Z;\ (\mathrm{proj}_{X}^{1}z)_{-1}=(\mathrm{proj}_{X}^{2}z)_{-1}\,\}.$
@@
Before proving Theorem A, we quote two results from the literature and prove a lemma. The first is a consequence of [49, Theorems 4.4 and 5.1] and the second is a result proved in [40].
Lemma 5.1 ([49, Lemma 3.1, Theorems 4.4 and 5.1]).
@@
We note that [40, Proposition 13.2] treats a slightly different setting, but the proof applies in our context because it uses only the finite-to-one coding properties (see also [9, Thm. 0.2] and [18, Thm. 10.5]).
Lemma 5.2 ([40, Proposition 13.2 and Step 2 of its proof]).
@@
Theorem 5.4 ([24, Chapter 5, Theorem 3.1.1]).

### v2

@@ 4. Measures of Relative Maximal Entropy on Topological Markov Shifts
We say that $\mu_{1}$ and $\mu_{2}$ are $\nu$-relatively orthogonal if
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, if $\mu_{1}$ and $\mu_{2}$ are ergodic measures of maximal $\nu$-relative entropy then $\mu_{1}$ and $\mu_{2}$ are $\nu$-relatively orthogonal.
@@
Let $(X,\sigma_{X})$ be an irreducible countable topological Markov shift with finite Gurevich topological entropy, $(Y,\sigma_{Y})$ an irreducible subshift of finite type and $\pi:X\to Y$ a surjective $1$-block factor map. Then, $(X,\sigma_{X})$ can only admit countably many ergodic measures of maximal $\nu$-relative entropy.
@@
Given an invariant measure $\mu$ of $(X,\sigma)$ such that $\pi_{*}\mu=\nu$, observe that $\mu_{n}:=(\mathrm{proj}_{X_{n}})_{*}\mu$ satisfies $(\pi_{n})_{*}\mu_{n}=\nu$. From the Kolmogorov-Sinai theorem (see [51, Theorem 4.9]) we have that
@@
where $\beta_{n}$ is the partition of $X_{n}$ 11-cylinders at coordinate 0. Since $\mathcal{B}(X)=\mathcal{B}\left(\bigcup_{n}\sigma(\alpha_{n})\right)$ from [51, Theorem 4.14]
@@ 4.1. Petersen-Quas-Shin construction
Now, we quote a result from [44] which is contained in the proof of Theorem 1 (see [44, Equation (45)]).
@@
We assume the notations of Theorem 4.6 and Lemma 4.8. Assume, for contradiction, that the theorem is false. Then there exist two ergodic measures of maximal $\nu$-relative entropy, $\mu^{1}$ and $\mu^{2}$, such that
@@
Define $\mu^{3}=(\mathrm{proj}^{3}_{X})_{*}\eta$ by the construction at the beginning of Section 4.1 (we keep the same notation, omitting the index $n$). We will show that
@@
$h_{\mu^{i}}(X)=\lim_{n\to\infty}h_{\mu^{i}_{n}}(X_{n})\ \text{for }i\in\{1,2,3\},\ \text{where}\ \mu_{n}^{i}:=(\mathrm{proj}_{X_{n}})_{*}\mu^{i}.$
@@
on an $\eta$-positive measure set contained in $S:=\{\,z\in Z;\ (\mathrm{proj}_{X}^{1}(z))_{-1}=(\mathrm{proj}_{X}^{2}(z))_{-1}\,\}.$
@@
Before proving Theorem A, we quote two results from the literature and prove a lemma. The first is a consequence of [54, Theorems 4.4 and 5.1] and the second is a result proved in [45].
Lemma 5.1 ([54, Lemma 3.1, Theorems 4.4 and 5.1]).
@@
We note that [45, Proposition 13.2] treats a slightly different setting, but the proof applies in our context because it uses only the finite-to-one coding properties (see also [10, Thm. 0.2] and [19, Thm. 10.5]).
Lemma 5.2 ([45, Proposition 13.2 and Step 2 of its proof]).
@@
Theorem 5.4 ([25, Chapter 5, Theorem 3.1.1]).